OpenStudy (anonymous):

Multiplication Property of Negative One Prove that -1x X= -X

OpenStudy (anonymous):

a.\[-1\times x +x=-1\times x+ 1\times x\] b.\[=(-1+1)(x)\] c.\[=0\times x\] d. =0 e. \[=-x+x\] f.\[\therefore -1 \times x + x = - x +x\] g.\[\therefore -1 \times x= -x\]

OpenStudy (anonymous):

could anyone help me w/ the reasons for this??? like mult. prop. of = or sth...